# Stabilizer Entanglement as a Magic Highway

### v2

@@ Appendix E Exact results
Appendix F Simulation method
In this section, we briefly discuss the numerical simulation method we used. We use two different methods for numerically evaluting the average $\overline{Y}$. The first one is a brute-force evaluation according to the definition of $Y$. However, this requires a direct enumeration of all $4^{N}$ Pauli strings, which can be time-consuming for larger system sizes. We thus adopt a second method, which is inspired by our analytical calculations to avoid summing over exponentially many Pauli strings. According to Eq. (17), there are only four different values for ${{\alpha}_{{{P}_{A}}}}({{a}_{k}},{{a}_{k}},{{a}_{k}},{{a}_{k}})$, depending on whether $P_{A}=I_{A}$ and $a_{k}=I_{A}$. Therefore, we only need to evaluate the corresponding $\alpha$’s to efficiently compute the average $\overline{Y}$. Similarly, there are four possible contributions from different $P$ in Eq. (20), so we can simply simulate these values to obtain an unbiased estimate for $\overline{Y}$.

### v3

@@ Appendix H Exact results
Appendix I Simulation method
In this section, we briefly discuss the numerical simulation method we used. We use two different methods for numerically evaluating the average $\overline{Y^{\rm lin}}$. The first one is a brute-force evaluation according to the definition of $Y^{\rm lin}$. However, this requires a direct enumeration of all $4^{N}$ Pauli strings, which can be time-consuming for larger system sizes. We thus adopt a second method, which is inspired by our analytical calculations to avoid summing over exponentially many Pauli strings. According to Eq. (21), there are only four different values for ${{\alpha}_{{{P}_{A}}}}({{a}_{k}},{{a}_{k}},{{a}_{k}},{{a}_{k}})$, depending on whether $P_{A}=I_{A}$ and $a_{k}=I_{A}$. Therefore, we only need to evaluate the corresponding $\alpha$’s to efficiently compute the average $\overline{Y^{\rm lin}}$. Similarly, there are four possible contributions from different $P$ in Eq. (24), so we can simply simulate these values to obtain an unbiased estimate for $\overline{Y^{\rm lin}}$.